water is taken from a well 11m deep in a bucket, the rope winding on a drum 35cm in diameter.through what angle does the handle turn in winding up 1 meter if rope? how many revolution of the handle does it take to bring the bucket up from the bottom? if the harm is 42cm long, how far does the hand of the winder travel when bringing the bucket up from the bottom
Answer by addingup(3677)

drum diameter = 35 cm = 0.35 m
radius = 0.35/2=0.175 m
height = 11 m
.
2 * pi * 0.175 = 1.0996. m
Number of revolutions = 11/1.0996 = 10 revolutions
.
then you say "if the harm is 42cm long" Sorry, I don't know what the harm is.